• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zhan

    Log entry time 00:02:23 on May18,2010

    Entry number 323029

    keyword=Shift Summary

    Swing shift summary:

    16:05: Shift starts with beam recovery, target vacummn keep decreasing.

    17:50: Ramping current up to 40uA, but we lost the control of left Q1, Jack called in.

    18:00: Called MCC, asked for control access

    18:10: Jack rebooted the IOC, left dipole alarm, back to normal in a couple of minutes.

    18:30: Jack fixed Q1, back to beam permit, ask for 5uA. Turned Septum magnets back on. Checked the spot++, ready for production.

    18:36: Ask for 50uA and ramping up slowly. Starting production!

    20:16: Compton HV tripped and reset.

    good production runs
    --------------------------
    run# 3918
    run# 3919
    run# 3920
    run# 3921
    run# 3922
    run# 3935
    run# 3936 continue
    ----------------------------